# Some arithmetic aspects of ortho-integral surfaces

### v1

@@ 2.4. Arithmetic Fuchsian groups coming from quadratic forms
where $\{\mathbf{e}_{1},\mathbf{e}_{2},\mathbf{e}_{3}\}$ is an orthogonal basis for $V$ such that $q({\bf e}_{1})<0$ In Section 5, we establish the connection between OI pair of pants and OI one-holed torus and a certain quadratic form that depends on a choice of an orthobasis. This connection allows a straightforward verification of arithmeticity of the double of these surfaces along their boundary components.
@@ 2.5. Hyperbolic identities
Figure 1. A pair of pants $P$ with the standard orthobasis =$\mathcal{H}=\{\alpha,\beta,\gamma\}$ and three geodesic boundary components $\delta_{\alpha},\delta_{\beta},$ and $\delta_{\gamma}$.
@@
  • Proof.
    It suffices to check the identity for $\operatorname{tr}_{\operatorname{SL}}^{2}(\delta_{\alpha})$ since the other two can be checked in a similar manner. Let $\eta$ be the embedded orthogeodesic that does not intersect $\alpha$ or $\beta$ and intersects $\gamma$ exactly once. Applying 2.3, we get
    $\cosh(\ell(\eta))=\frac{2(a^{2}+b^{2}+c^{2}+2abc)}{c^{2}-1}-1.$
    The orthogeodesics $\beta$ and $\eta$ bound a right-angled hexagons with a triple alternating sides $(\beta$, $\beta,\eta)$. In this right-angled hexagon, the boundary $\delta_{\alpha}$ is opposite to the sides $\eta$. Applying Equation 5, we get
    $\cosh(\delta_{\alpha})=\frac{b^{2}+\cosh(\eta)}{b^{2}-1}=\frac{2(bc+a)^{2}}{(b^{2}-1)(c^{2}-1)}-1=2\cosh(\delta_{\alpha}/2)^{2}-1,$
    which implies that
    $\operatorname{tr}^{2}_{\operatorname{SL}}(\delta_{\alpha})=4\cosh^{2}(\delta_{\alpha}/2)=\frac{4(bc+a)^{2}}{(b^{2}-1)(c^{2}-1)}.$
@@
Let $\vec{\eta}_{k}$ be the oriented orthogeodesic with cutting sequence of length $12k+$ 1 of the form $(\gamma,\beta,\dots,\gamma)$ where $k\geq 0$. The cosh-length of the orthogeodesic $\eta_{k}$ can be computed recursively by
@@ 3. The finiteness of OI surfaces
A hyperbolic surface $X$ is $d$-almost-ortho-integral ($d$-AOI) for some $d\in\mathbb{Z}^{+}$ if all orthogeodesics of $X$ have cosh-lengths in $\frac{1}{d}\mathbb{Z}$.
@@
Let $S$ be a hyperbolic surface with nonempty totally geodesic boundary. Fix a boundary component $\delta$ of $S$. We say that $S$ is partially $d$-AOI with respect to $\delta$ if the cosh-lengths of all orthogeodesics with both endpoints on $\delta$ lies in $\frac{1}{d}\mathbb{Z}$.
@@
Let $P$ be a hyperbolic pair of pants with a standard orthobasis =$\mathcal{H}=\{\alpha,\beta,\gamma\}$. Denote $(a,b,c):=(\cosh\alpha,\cosh\beta,\cosh\gamma)$. If $P$ is partially $d$-AOI with respect to the boundary $\delta_{\gamma}$, then
(8) $\operatorname{tr}_{\operatorname{SL}}^{2}(\alpha)=\frac{4(bc+a)^{2}}{(b^{2}-1)(c^{2}-1)},\quad\text{and}\quad\operatorname{tr}_{\operatorname{SL}}^{2}(\beta)=\frac{4(ca+b)^{2}}{(c^{2}-1)(a^{2}-1)}$
@@
  • Proof.
    It suffices to prove that $\operatorname{tr}_{\operatorname{SL}}^{2}(\alpha)\in\mathbb{N}$. The other case followed by a similar argument. Let $\vec{\eta}_{k}$ be the oriented orthogeodesic with cutting sequence of length $12k+$ 1 of the form $(\gamma,\beta,\dots,\gamma)$ where $k\geq 0$. By 2.8, we have $\cosh(\eta_{0})=s-1$, $\cosh(\eta_{1})=\tau s-1$ and $\cosh(\eta_{2})=\tau^{2}s-2\tau s+s-1$ where
    (9) $\tau=\operatorname{tr}_{\operatorname{SL}}^{2}(\alpha)=\frac{4(bc+a)^{2}}{(b^{2}-1)(c^{2}-1)}\quad\text{and}\quad s=\frac{2(a^{2}+b^{2}+c^{2}+2abc-1)}{c^{2}-1}.$
    Since $P$ is partially $d$-AOI with respect to the boundary $\delta_{\gamma}$, $\cosh(\eta_{k})\in\frac{1}{d}\mathbb{Z}$ for all $k\in\mathbb{N}$. The initial conditions for $\cosh(\eta_{k})$ implies that $\{s,\tau s,\tau^{2}s\}\subset\frac{1}{d}\mathbb{Z}$. From the initial conditions and the recursive formula for $\cosh(\eta_{k})$ in Equation 7, the formula for $\cosh(\eta_{k})$ is a $\mathbb{Z}$-linear combination of $\{\tau^{k}s,\tau^{k-1}s\dots,s,1\}$ where the coefficient of $\tau^{k}s$ is one. Arguing by induction, we get that $\tau^{k}s\in\frac{1}{d}\mathbb{Z}$ for all integers $k\geq 0$. We see that $\tau\in\mathbb{Q}$ since $s,\tau s\in\frac{1}{d}\mathbb{Z}$. Thus, $\tau^{k}s\in\frac{1}{d}\mathbb{Z}$ for all integers $k\geq 0$ if and only if $\tau=\operatorname{tr}_{\operatorname{SL}}^{2}(\alpha)\in\mathbb{N}$. ∎
The following proposition provides the key bounds needed to establish the finiteness of the set of $d$–AOI surfaces. The crucial observation is that for any two (not necessary distinct) closed geodesics on a hyperbolic surface $X$, there exists an immersed pair of pants bounded by these two geodesics. This pair of pants can be constructed explicitly as follows. Let $\Gamma\leq\operatorname{PSL}_{2}(\mathbb{R})$ be the image of $\pi_{1}(X)$ under the holonomy representation. For any closed geodesics $\mu_{1}$ and $\mu_{2}$ on $X$, we choose two disjoint lifts $\widetilde{\mu}_{1}$ of $\mu_{1}$ and $\widetilde{\mu}_{2}$ of $\mu_{2}$ in the universal cover $\widetilde{X}\subset\mathbb{H}^{2}$. The isometries $g_{\mu_{1}}\in\Gamma$ stabilizing $\widetilde{\mu}_{1}$ and $g_{\mu_{2}}\in\Gamma$ stabilizing $\widetilde{\mu}_{2}$ generate a subgroup $\Delta$ in $\Gamma$. Let $C(\Lambda_{\Delta})$ be the convex hull of the limit set of $\Delta$. Then the quotient $C(\Lambda_{\Delta})/\Delta$ is isometric to a pair of pants $P$. The natural inclusion of $C(\Lambda_{\Delta})$ in $\widetilde{X}\subset\mathbb{H}^{2}$ induces an isometric immersion of the pair of pants $P$ into $X$.
@@
Immersed pants were used in [BPT25, Par20] to extend Basmajian’s identity beyond hyperbolic surfaces with geodesic boundary.
@@
Let $X\in\mathcal{M}_{g,n}$ be a $d$-AOI surface of genus $g$ and $n$ geodesic boundary components. Let $L$ be the total boundary length of $S$. Then
@@
  • Proof.
    $(i)$ Let $P$ be an immersed pair of pants on $X$ with three geodesic boundary components $\delta_{\alpha},\delta_{\beta},\delta_{\gamma},$ where $\delta_{\alpha}=\mu$ and $\delta_{\gamma}\in\partial X$. The pair of pants $P$ is constructed as described in the preceding paragraph. Since $X$ is a $d$–AOI surface, $P$ is partially $d$–AOI with respect to $\delta_{\gamma}$. By 3.4, $\operatorname{tr}_{\operatorname{SL}}^{2}(\mu)\in\mathbb{Z}$.
    $(ii)$ Since $X$ is $d$–AOI, $\cosh(\operatorname{osys}(X))$ is greater than or equal to $1+\frac{1}{d}$. Applying 2.5, we have
@@
  • Proof of 3.2.
    3.6 provides uniform upper and lower bounds on the length of each boundary component of $X$ and establishes a lower bound on the systole of $X$. By Mumford’s compactness criterion 2.1, the set of $d$–AOI surfaces $X\in\mathcal{M}_{g,n}$ is contained within a compact subset of the moduli space $\mathcal{M}_{g,n}$.
    Since the cosh-lengths of orthogeodesics of $d$–AOI surfaces can only take value in a fixed discrete subset of $\mathbb{R}$, the set of $d$–AOI surfaces of genus $g$ and $n$ boundary components is discrete in the moduli space $\mathcal{M}_{g,n}$. Together with the compactness, we conclude that there are only finitely many $d$–AOI surfaces for a fixed $d$ and fixed topological complexity.
@@
The discreteness in the moduli space $\mathcal{M}_{g,n}$ still holds for the set of partially $d$–AOI surfaces of genus $g$ with $n$ boundary components since $\operatorname{tr}_{\operatorname{SL}}^{2}(\mu)\in\mathbb{Z}$ for every closed geodesic $\mu$ on these surfaces. However, finiteness may fail for this broader class, as we do not have a uniform lower bound on the orthosystole, which in turn prevents us from deriving a uniform upper bound on the total boundary length.
4. Classification of OI pants and OI one-holed tori
4.1. Classification of OI pants
The following result provides another way to prove the finiteness of ortho-integral pants without using the Mumford’s compactness criterion and offers a complete classification of them.
@@ 4.2. Classification of OI one-holed tori
    $22\cosh(L/2)=2\frac{a^{2}+q}{a^{2}-1}=2\frac{a^{2}-b+\frac{a+c}{b-1}\left(1+\frac{(a+b)^{2}}{c-1}+a\right)}{a^{2}-1}=\frac{2(a+b+c-1)^{2}}{(a-1)(b-1)(c-1)}+$ 2 .
@@
      This would force $a<-1$, which is a contradiction.
@@
    $2(\cosh(L/2)-1)(a-1)=\frac{2(a+b+c-1)^{2}}{(b-1)(c-1)}\geq\frac{2((c-b-1)+b+c-1)^{2}}{(b-1)(c-1)}=\frac{8(c-1)}{b-1}\geq 8$
    The equality happens if and only if $a=c-b-1$ and $b=c$, which implies that $a=-1$, a contradiction. Rearranging the inequality above, we obtain a strict lower bound
@@
    Therefore $a,b,c,|\operatorname{tr}_{\operatorname{SL}}(\partial T)|\in\mathbb{N}$ satisfying that $2\leq a\leq b\leq c\leq 107$, $c-b-1\leq a\leq 20$ and $3\leq|\operatorname{tr}_{\operatorname{SL}}(\partial T)|\leq 52$
@@ 5. Ortho-integral surfaces and quadratic forms
In this section, we describe some connections between OI one-holed tori and OI pants with quadratic forms. We leverage these connections to construct examples of infinitely many pairwise non-commensurable OI surfaces and to compute arithmetic invariants of a one-holed torus and a pair of pants, from the data of a decomposition of the surface into isometric right-angled hexagons.
@@ 5.1. Connection with quadratic form
where $q({\bf u}_{*})>0$, $q({\bf v}_{*})<0$, $q({\bf w}_{*})<0$ and $q({\bf e}_{i})>0$ for all $*\in\{\alpha,\beta,\gamma\}$ and $1\leq i\leq 3$. Therefore, the quadratic form $q_{(a,b,c)}$ has signature (+,+,)(+,+,-)( + , + , - ) for all real numbers $a,b,c>1$. Since $q({\bf v}_{\alpha})<0$, we can choose a model for the hyperbolic plane by projectivizing the component of $q({\bf x})<0$ that contains ${\bf v}_{\alpha}$. The hyperbolic plane is modeled on the projectivization $P(\mathcal{C}^{+})$ of the set
@@
The orthogonal complement ${\bf u}_{*}^{\perp}\subset\mathbb{R}^{3}$ projects to an (infinite) geodesic in $P(\mathcal{C}^{+})$ that contains the geodesic segment $[{\bf v}_{*},{\bf w}_{*}]$ for each $*\in\{\alpha,\beta,\gamma\}$. The orthogonal complement ${\bf e}_{i}^{\perp}\subset\mathbb{R}^{3}$ projects to an (infinite) geodesic in $P(\mathcal{C}^{+})$ that contains the geodesic segment * for each $*\in\{\alpha,\beta,\gamma\}$.
  • Proof.
    Since $q({\bf v}_{*})<0$ and $q({\bf w}_{*})<0$ for any $*\in\{\alpha,\beta,\gamma\}$, the vectors ${\bf v}_{*}$ and ${\bf w}_{*}$ project to six points on $P(\mathcal{C}^{+})$. Since $q({\bf u}_{*})>0$ and $q({\bf e}_{i})>0$ for any $*\in\{\alpha,\beta,\gamma\}$ and $1\leq i\leq 3$, the orthogonal complement ${\bf u}_{*}^{\perp}$ and ${\bf e}_{i}^{\perp}$ projects to geodesics on $P(\mathcal{C}^{+})$. The fact that each set in Equation 20 is an orthogonal basis for $(V,q)$ implies that the convex hull of $\{{\bf v}_{\alpha},{\bf v}_{\beta},{\bf v}_{\gamma},{\bf w}_{\alpha},{\bf w}_{\beta},{\bf w}_{\gamma}\}$ is an all-right angled hexagon and that the orthogonal complements ${\bf u}_{*}^{\perp}$ and ${\bf e}_{i}^{\perp}$ projects to the sides of the hexagon as claimed. Using Equation 21, we compute the cosh of the distance between the points ${\bf v}_{*}$ and ${\bf w}_{*}$ to be
@@
Naturally, for $*\in\{\alpha,\beta,\gamma\}$ the matrices $f_{*}$ and $g_{*}$ from 2.3 also have geometric interpretation. In particular, $f_{*}$ and $g_{*}$ preserve the quadratic form $q$ and $\mathcal{C}^{+}$, thus define non-trivial isometries of the hyperboloid model. Since $f_{\alpha}$ fixes ${\bf v}_{\beta}$, ${\bf w}_{\gamma}$, the isometry $f_{\alpha}$ is a reflection across $\alpha$. Similarly, $f_{\beta}$ is a reflection across $\beta$, and $f_{\gamma}$ is a reflection across $\gamma$. The matrix $g_{*}$ has order 2 and $\det(g_{*})=1$ for any $*\in\{\alpha,\beta,\gamma\}$, and thus defines an order two rotation. Since $g_{\alpha}({\bf v}_{\beta})={\bf w}_{\gamma}$, $g_{\alpha}$ is an order two rotation about the midpoint of $\alpha$. Similarly, $g_{\beta}$ and $g_{\gamma}$ are order two rotations about the midpoints of $\beta$ and $\gamma$ respectively. The following proposition is immediate from the discussion above.
@@
Let $P(a,b,c)$ be a pair of pants with a standard orthobasis $\mathcal{H}_{P}=\{\alpha,\beta,\gamma\}$ and $(a,b,c):=(\cosh\alpha,\cosh\beta,\cosh\gamma)$. The group $G_{P}$ generated by $f_{\alpha}$, $f_{\beta}$ and $f_{\gamma}$ is a discrete subgroup of $\operatorname{O}^{+}(q_{a,b,c};\mathbb{R})$. The universal cover $\widetilde{P}(a,b,c)$ of $P(a,b,c)$ is isometric to a tiling of a convex subset of $P(\mathcal{C}^{+})$ by the orbit of $H(a,b,c)$ under the action of $G_{P}$. The subgroup $\Gamma_{P}:=\langle f_{\alpha}f_{\beta},f_{\alpha}f_{\gamma}\rangle$ has index two in $G_{P}$, and $\widetilde{P}(a,b,c)/\Gamma_{P}$ is isometric to $P(a,b,c)$.
Let $T(a,b,c)$ be a one-holed torus with a standard orthobasis $\mathcal{H}_{T}=\{\alpha,\beta,\gamma\}$ and $(a,b,c):=(\cosh\alpha,\cosh\beta,\cosh\gamma)$. The group $G_{T}$ generated by $g_{\alpha}$, $g_{\beta}$ and $g_{\gamma}$ is a discrete subgroup of $\operatorname{O}^{+}(q_{a,b,c};\mathbb{R})$. The universal cover $\widetilde{T}(a,b,c)$ of $T(a,b,c)$ is isometric to a tiling of a convex subset of $P(\mathcal{C}^{+})$ by the orbit of $H(a,b,c)$ under the action of $G_{T}$. The subgroup $\Gamma_{T}:=\langle g_{\alpha}g_{\beta},g_{\alpha}g_{\gamma}\rangle$ has index two in $G_{T}$, and $\widetilde{T}(a,b,c)/\Gamma_{T}$ is isometric to $T(a,b,c)$.
5.2. Verifying arithmeticity of the double of OI pants and OI one-holed torus
In this section, we compute the arithmetic invariants of OI one-holed torus and OI pants. We verify that the doubles of OI one-holed tori and OI pants are all arithmetic surfaces of genus 2. The following theorem provides a criterion for determining whether a finite-covolume Fuchsian group is arithmetic.

### v2

@@ 2.4. Arithmetic Fuchsian groups coming from quadratic forms
where $\{\mathbf{e}_{1},\mathbf{e}_{2},\mathbf{e}_{3}\}$ is an orthogonal basis for $V$ such that $q({\bf e}_{1})<0$ . In Section 5, we establish the connection between OI pair of pants and OI one-holed torus and a certain quadratic form that depends on a choice of an orthobasis. This connection allows a straightforward verification of arithmeticity of the double of these surfaces along their boundary components.
@@ 2.5. Hyperbolic identities
Figure 2. A pair of pants $P$ with the standard orthobasis $\mathcal{H}=\{\alpha,\beta,\gamma\}$ and three geodesic boundary components $\delta_{\alpha},\delta_{\beta},$ and $\delta_{\gamma}$.
@@
It suffices to check the identity for $\operatorname{tr}_{\operatorname{SL}}^{2}(\delta_{\alpha})$ since the other two can be checked in a similar manner. Applying Equation 5 to the hexagon with alternating edges $\alpha$, $\beta$ and $\gamma$, we get
$\cosh\left(\frac{\ell(\delta_{\alpha})}{2}\right)=\frac{(bc+a)}{\sqrt{(b^{2}-1)(c^{2}-1)}}.$
The lemma follows since $\operatorname{tr}^{2}_{\operatorname{SL}}(\delta_{\alpha})=4\cosh^{2}\left(\frac{\ell(\delta_{\alpha})}{2}\right)$.
@@
Let $\eta_{k}$ be the oriented orthogeodesic with cutting sequence of length $2k+1$ of the form $(\gamma,\beta,\dots,\gamma)$ where $k\geq 0$. The cosh-length of the orthogeodesic $\eta_{k}$ can be computed recursively by
@@ 3. The finiteness of OI surfaces
A hyperbolic surface $X$ is $d$-almost-ortho-integral ($d$-AOI) for some $d\in\mathbb{Z}^{+}$ if the cosh of the length of every orthogeodesic of $X$ is a number in $\frac{1}{d}\mathbb{Z}$.
@@
Let $X$ be a hyperbolic surface with nonempty totally geodesic boundary. Fix a boundary component $\delta$ of $X$. We say that $X$ is partially $d$-AOI with respect to $\delta$ if the cosh of the length of every orthogeodesic with both endpoints on $\delta$ lies in $\frac{1}{d}\mathbb{Z}$.
@@
Let $P$ be a hyperbolic pair of pants with a standard orthobasis $\mathcal{H}=\{\alpha,\beta,\gamma\}$. Denote $(a,b,c):=(\cosh\ell(\alpha),\cosh\ell(\beta),\cosh\ell(\gamma))$. If $P$ is partially $d$-AOI with respect to the boundary $\delta_{\gamma}$, then
(8) $\operatorname{tr}_{\operatorname{SL}}^{2}(\delta_{\alpha})=\frac{4(bc+a)^{2}}{(b^{2}-1)(c^{2}-1)},\quad\text{and}\quad\operatorname{tr}_{\operatorname{SL}}^{2}(\delta_{\beta})=\frac{4(ca+b)^{2}}{(c^{2}-1)(a^{2}-1)}$
@@
It suffices to prove that $\operatorname{tr}_{\operatorname{SL}}^{2}(\delta_{\alpha})\in\mathbb{N}$. The other case followed by a similar argument. Let $\eta_{k}$ be the oriented orthogeodesic with cutting sequence of length $2k+1$ of the form $(\gamma,\beta,\dots,\gamma)$ where $k\geq 0$. By 2.8, we have $\cosh(\ell(\eta_{0}))=s-1$, $\cosh(\ell(\eta_{1}))=\tau s-1$ and $\cosh(\ell(\eta_{2}))=\tau^{2}s-2\tau s+s-1$ where
(9) $\tau=\operatorname{tr}_{\operatorname{SL}}^{2}(\delta_{\alpha})=\frac{4(bc+a)^{2}}{(b^{2}-1)(c^{2}-1)}\quad\text{and}\quad s=\frac{2(a^{2}+b^{2}+c^{2}+2abc-1)}{c^{2}-1}.$
Since $P$ is partially $d$-AOI with respect to the boundary $\delta_{\gamma}$, $\cosh(\ell(\eta_{k}))\in\frac{1}{d}\mathbb{Z}$ for all $k\in\mathbb{N}$. The initial conditions for $\cosh(\ell(\eta_{k}))$ implies that $\{s,\tau s,\tau^{2}s\}\subset\frac{1}{d}\mathbb{Z}$. From the initial conditions and the recursive formula for $\cosh(\ell(\eta_{k}))$ in Equation 7, the formula for $\cosh(\ell(\eta_{k}))$ is a $\mathbb{Z}$-linear combination of $\{\tau^{k}s,\tau^{k-1}s\dots,s,1\}$ where the coefficient of $\tau^{k}s$ is one. Arguing by induction, we get that $\tau^{k}s\in\frac{1}{d}\mathbb{Z}$ for all integers $k\geq 0$. We see that $\tau\in\mathbb{Q}$ since $s,\tau s\in\frac{1}{d}\mathbb{Z}$. Thus, $\tau^{k}s\in\frac{1}{d}\mathbb{Z}$ for all integers $k\geq 0$ if and only if $\tau=\operatorname{tr}_{\operatorname{SL}}^{2}(\delta_{\alpha})\in\mathbb{N}$. ∎
The following proposition provides the key bounds needed to establish the finiteness of the set of dd–AOI surfaces. The crucial observation is that for any two (not necessary distinct) closed geodesics on a hyperbolic surface $X$, there exists an immersed pair of pants bounded by these two geodesics. This pair of pants can be constructed explicitly as follows. Let $\Gamma\leq\operatorname{PSL}_{2}(\mathbb{R})$ be the image of $\pi_{1}(X)$ under the holonomy representation. For any closed geodesics $\mu_{1}$ and $\mu_{2}$ on $X$, we choose two disjoint lifts $\widetilde{\mu}_{1}$ of $\mu_{1}$ and $\widetilde{\mu}_{2}$ of $\mu_{2}$ in the universal cover $\widetilde{X}\subset\mathbb{H}^{2}$. The isometries $g_{\mu_{1}}\in\Gamma$ stabilizing $\widetilde{\mu}_{1}$ and $g_{\mu_{2}}\in\Gamma$ stabilizing $\widetilde{\mu}_{2}$ generate a subgroup $\Delta$ in $\Gamma$. Let $C(\Lambda_{\Delta})$ be the convex hull of the limit set of $\Delta$. Then the quotient $C(\Lambda_{\Delta})/\Delta$ is isometric to a pair of pants $P$. The natural inclusion of $C(\Lambda_{\Delta})$ in $\widetilde{X}\subset\mathbb{H}^{2}$ induces an isometric immersion of the pair of pants $P$ into $X$.
@@
Immersed pairs of pants were used in [BPT25, Par20] to extend Basmajian’s identity beyond hyperbolic surfaces with geodesic boundary.
@@
Let $X\in\mathcal{M}_{g,n}$ be a $d$-AOI surface. Let $L$ be the total boundary length of $X$. Then
@@
$(i)$ Let $P$ be an immersed pair of pants on $X$ with three geodesic boundary components $\delta_{\alpha},\delta_{\beta},\delta_{\gamma},$ where $\delta_{\alpha}=\mu$ and $\delta_{\gamma}\in\partial X$. The pair of pants $P$ is constructed as described in the preceding paragraph. Since $X$ is a dd–AOI surface, $P$ is partially dd–AOI with respect to $\delta_{\gamma}$. By 3.4, $\operatorname{tr}_{\operatorname{SL}}^{2}(\mu)\in\mathbb{Z}$.
$(ii)$ Since $X$ is dd–AOI, $\cosh(\operatorname{osys}(X))$ is greater than or equal to $1+\frac{1}{d}$. Applying 2.5, we have
@@
3.6 provides uniform upper and lower bounds on the length of each boundary component of $X$ and establishes a lower bound on the systole of $X$. By Mumford’s compactness criterion 2.1, the set of dd–AOI surfaces $X\in\mathcal{M}_{g,n}$ is contained within a compact subset of the moduli space $\mathcal{M}_{g,n}$.
Since the cosh-lengths of orthogeodesics of dd–AOI surfaces can only take value in a fixed discrete subset of $\mathbb{R}$, the set of dd–AOI surfaces is discrete in the moduli space $\mathcal{M}_{g,n}$. Together with the compactness, we conclude that there are only finitely many dd–AOI surfaces for a fixed $d$ and fixed topological complexity.
@@
The discreteness in the moduli space $\mathcal{M}_{g,n}$ still holds for the set of partially dd–AOI surfaces since $\operatorname{tr}_{\operatorname{SL}}^{2}(\mu)\in\mathbb{Z}$ for every closed geodesic $\mu$ on these surfaces. However, finiteness may fail for this broader class, as we do not have a uniform lower bound on the orthosystole, which in turn prevents us from deriving a uniform upper bound on the total boundary length.
4. Classification of OI pairs of pants and OI one-holed tori
4.1. Classification of OI pairs of pants
The following result provides another way to prove the finiteness of ortho-integral pairs of pants without using the Mumford’s compactness criterion and offers a complete classification of them.
@@ 4.2. Classification of OI one-holed tori
$2\cosh(L/2)=2\frac{a^{2}+q}{a^{2}-1}=2\frac{a^{2}-b+\frac{a+c}{b-1}\left(1+\frac{(a+b)^{2}}{c-1}+a\right)}{a^{2}-1}=\frac{2(a+b+c-1)^{2}}{(a-1)(b-1)(c-1)}+2.$
@@
    This would force a<1a<-1, which is a contradiction.
@@
$2(\cosh(L/2)-1)(a-1)=\frac{2(a+b+c-1)^{2}}{(b-1)(c-1)}\geq\frac{2((c-b-1)+b+c-1)^{2}}{(b-1)(c-1)}=\frac{8(c-1)}{b-1}\geq 8.$
The equality happens if and only if a=cb1a=c-b-1 and $b=c$, which implies that a=1a=-1, a contradiction. Rearranging the inequality above, we obtain a strict lower bound
@@
Therefore $a,b,c,|\operatorname{tr}_{\operatorname{SL}}(\partial T)|\in\mathbb{N}$ satisfying that $2\leq a\leq b\leq c\leq 107$, $c-b-1\leq a\leq 20$ and $3\leq|\operatorname{tr}_{\operatorname{SL}}(\partial T)|\leq 52.$
@@ 5. Ortho-integral surfaces and quadratic forms
In this section, we describe some connections between OI one-holed tori and OI pairs of pants with quadratic forms. We leverage these connections to construct examples of infinitely many pairwise non-commensurable OI surfaces and to compute arithmetic invariants of a one-holed torus and a pair of pants, from the data of a decomposition of the surface into isometric right-angled hexagons.
@@ 5.1. Connection with quadratic form
where $q({\bf u}_{*})>0$, $q({\bf v}_{*})<0$, $q({\bf w}_{*})<0$ and $q({\bf e}_{i})>0$ for all $\in\{\alpha,\beta,\gamma\}$ and $1\leq i\leq 3$. Therefore, the quadratic form $q_{(a,b,c)}$ has signature (+,+,)(+,+,-) for all real numbers $a,b,c>1$. Since $q({\bf v}_{\alpha})<0$, we can choose a model for the hyperbolic plane by projectivizing the component of $q({\bf x})<0$ that contains ${\bf v}_{\alpha}$. The hyperbolic plane is modeled on the projectivization $P(\mathcal{C}^{+})$ of the set
@@
The orthogonal complement ${\bf u}_{*}^{\perp}\subset\mathbb{R}^{3}$ projects to an (infinite) geodesic in $P(\mathcal{C}^{+})$ that contains the geodesic segment ${\bf v}_{*},{\bf w}_{*}]$ for each $\in\{\alpha,\beta,\gamma\}$. The orthogonal complement ${\bf e}_{i}^{\perp}\subset\mathbb{R}^{3}$ projects to an (infinite) geodesic in $P(\mathcal{C}^{+})$ that contains the geodesic segment * for each $\in\{\alpha,\beta,\gamma\}$.
Since $q({\bf v}_{*})<0$ and $q({\bf w}_{*})<0$ for any $\in\{\alpha,\beta,\gamma\}$, the vectors $\bf v$$_{*}$ and $\bf w$$_{*}$ project to six points on $P(\mathcal{C}^{+})$. Since $q({\bf u}_{*})>0$ and $q({\bf e}_{i})>0$ for any $\in\{\alpha,\beta,\gamma\}$ and $1\leq i\leq 3$, the orthogonal complement $\bf u$$_{*}^{\perp}$ and ${\bf e}_{i}^{\perp}$ projects to geodesics on $P(\mathcal{C}^{+})$. The fact that each set in Equation 20 is an orthogonal basis for $(V,q)$ implies that the convex hull of $\{{\bf v}_{\alpha},{\bf v}_{\beta},{\bf v}_{\gamma},{\bf w}_{\alpha},{\bf w}_{\beta},{\bf w}_{\gamma}\}$ is an all-right angled hexagon and that the orthogonal complements $\bf u$$_{*}^{\perp}$ and ${\bf e}_{i}^{\perp}$ projects to the sides of the hexagon as claimed. Using Equation 21, we compute the cosh of the distance between the points $\bf v$$_{*}$ and $\bf w$$_{*}$ to be
@@
Naturally, for $\in\{\alpha,\beta,\gamma\}$ the matrices $f_{*}$ and $g_{*}$ from 2.3 also have a geometric interpretation. In particular, $f_{*}$ and $g_{*}$ preserve the quadratic form $q$ and $\mathcal{C}^{+}$, thus defining non-trivial isometries of the hyperboloid model. Since $f_{\alpha}$ fixes ${\bf v}_{\beta}$, ${\bf w}_{\gamma}$, the isometry $f_{\alpha}$ is a reflection across $\alpha$. Similarly, $f_{\beta}$ is a reflection across $\beta$, and $f_{\gamma}$ is a reflection across $\gamma$. The matrix $g_{*}$ has order 2 and $\det(g_{*})=1$ for any $\in\{\alpha,\beta,\gamma\}$, and thus defines an order two rotation. Since $g_{\alpha}({\bf v}_{\beta})={\bf w}_{\gamma}$, $g_{\alpha}$ is an order two rotation about the midpoint of $\alpha$. Similarly, $g_{\beta}$ and $g_{\gamma}$ are order two rotations about the midpoints of $\beta$ and $\gamma$ respectively. The following proposition is an immediate consequence of the discussion above. The key idea of 5.2 was communicated to us by Greg McShane. We would like to thank him for sharing his insights.
@@
Let $P(a,b,c)$ be a pair of pants with a standard orthobasis $\mathcal{H}_{P}=\{\alpha,\beta,\gamma\}$ and $(a,b,c):=(\cosh\ell(\alpha),\cosh\ell(\beta),\cosh\ell(\gamma))$. The group $G_{P}$ generated by $f_{\alpha}$, $f_{\beta}$ and $f_{\gamma}$ is a discrete subgroup of $\operatorname{O}^{+}(q_{a,b,c};\mathbb{R})$. The universal cover $\widetilde{P}(a,b,c)$ of $P(a,b,c)$ is isometric to a tiling of a convex subset of $P(\mathcal{C}^{+})$ by the orbit of $H(a,b,c)$ under the action of $G_{P}$. The subgroup $\Gamma_{P}:=\langle f_{\alpha}f_{\beta},f_{\alpha}f_{\gamma}\rangle$ has index two in $G_{P}$, and $\widetilde{P}(a,b,c)/\Gamma_{P}$ is isometric to $P(a,b,c)$.
Let $T(a,b,c)$ be a one-holed torus with a standard orthobasis $\mathcal{H}_{T}=\{\alpha,\beta,\gamma\}$ and $(a,b,c):=(\cosh\ell(\alpha),\cosh\ell(\beta),\cosh\ell(\gamma))$. The group $G_{T}$ generated by $g_{\alpha}$, $g_{\beta}$ and $g_{\gamma}$ is a discrete subgroup of $\operatorname{O}^{+}(q_{a,b,c};\mathbb{R})$. The universal cover $\widetilde{T}(a,b,c)$ of $T(a,b,c)$ is isometric to a tiling of a convex subset of $P(\mathcal{C}^{+})$ by the orbit of $H(a,b,c)$ under the action of $G_{T}$. The subgroup $\Gamma_{T}:=\langle g_{\alpha}g_{\beta},g_{\alpha}g_{\gamma}\rangle$ has index two in $G_{T}$, and $\widetilde{T}(a,b,c)/\Gamma_{T}$ is isometric to $T(a,b,c)$.
5.2. Verifying arithmeticity of the double of OI pair of pants and OI one-holed torus
In this section, we compute the arithmetic invariants of OI one-holed torus and OI pair of pants. We verify that the doubles of OI one-holed tori and OI pairs of pants are all arithmetic surfaces of genus 2. The following theorem provides a criterion for determining whether a finite-covolume Fuchsian group is arithmetic.